Road Warrior Corporation began operations early in the current year, building luxury motor homes. During the year, the company started and completed 50 motor homes at a cost of $60,000 per unit. Of these, 48 were sold for $100,000 each and two remain in finished goods inventory. In addition, the company had six partially completed units in its factory at year-end. Total costs for the year (summarized alphabetically) were as follows. Direct materials used $ 800,000 Direct labor 990,000 Income tax expense 100,000 General and administrative expenses 500,000 Manufacturing overhead 1,650,000 Selling expenses 500,000

Required:

a. Compute the total manufacturing costs charged to work in process for the current year.

b. Compute the cost of finished goods manufactured for the current year.

c. Compute the cost of goods sold for the current year.

d. Compute the gross profit on sales for the current year.

e. Compute the ending inventories of (1) work in process and (2) finished goods for the current year.
Answer #1

Cost of finished goods manufactured for the current year

Cost per unit of finished good manufactured * Number of finished goods units manufactured = Cost of finished goods manufactured

Cost of finished goods manufactured = 60000 * 50 = $3,000,000

Calculation of total manufacturing costs charged to work in progress for the current year

Total manufacturing costs = Direct materials + Direct labor + Manufacturig overhead = 800000 + 990000 + 1650000 = $3,440,000

Total manufacturing costs charged to work in progress = Total manufacturing costs - Cost of finished goods manufactured

Total manufacturing costs charged to work in progress = 3440000 - 3000000 = $440,000

Calculation of cost of goods sold for the current year

Manufacturing cost of goods sold = Cost per unit of finished good manufactured * Number of units sold = 60000 * 48 = $2,880,000

General and administrative expenses attributable to the goods sold = (Manufacturing cost of goods sold/Total manufacturing cost) * General and administrative expenses = (2880000/3440000)*500000 = $418,605

Cost of goods sold = Manufacturing cost of goods sold + General and administrative expenses attributable to the goods sold

Cost of goods sold = 2880000 + 418605 = $32,98,605

Calculation of gross profit on sales for the current year

Gross profit = Total sales - Manufacturing cost of goods sold = (100000*48) - (60000*48) = $1,920,000

Gross profit on sales = Gross profit/Sales = 1920000/4800000 = 40%

Computation of ending inventories of

1.) Work in Progress:

General and administrative expenses attributable to the work in progress = (Total manufacturing costs charged to work in progress/Total manufacturing cost) * General and administrative expenses = (440000/3440000)*500000 = $63,953

Closing inventory of work in progress = Total manufacturing costs charged to work in progress + General and administrative expenses attributable to the work in progress = 440000 + 63953 = $503,953

2.) Finished goods:

General and administrative expenses attributable to the finished goods = (Total manufacturing costs charged to finished goods/Total manufacturing cost) * General and administrative expenses = [(60000*2)/3440000]*500000 = $17,442

Closing inventory of Finished Goods = Total manufacturing costs charged to finished goods + General and administrative expenses attributable to finished goods = (60000*2) + 63953 = $183,953